Based on this Histogram, what percentage of data are below 41? 8 6 2 1 11-20 21-30 31-40 41-50 51-60 O 40% o 50% 65% 75%
Question 9 1 pts Calculate sample variance for the following data collection: 10, 25, 28, 35 (when you calculate variance, divide by N-1 not by O 85 O 150 O 180
Answer #1

Based on histogram what percentage of data are below 41?

- Percentage of data below 41

= (Count (11-20) + Count (21-30) + Count (31-40) ) / total * 100 %

= (1+4+7) / (1+4+7+3+1) * 100 %

= 0.75 * 100%

= 75%

9)

Sample: 10,25,28,35

Sample mean = (10+25+28+35)/4 = 24.5

using N-1, we get

sample variance = {(10-24.5)^2 + (25-24.5)^2 + (28-24.5)^2 + (35-24.5)^2}/ (4-1)

= 333/3

= 111